### v2

# Biharmonic nonlinear vector field equations in dimension 44

Abstract
Following the approach of Brezis and Lieb [8], we prove the existence of a ground state solution for the biharmonic nonlinear vector field equations in the limiting case of space dimension $4$. Our results complete those obtained by Mederski and Siemianowski [16] for dimensions $d\geq 5$. We also extend the biharmonic logarithmic Sobolev inequality in [16] to dimension $4$.
keywords:
biharmonic , ground state , elliptic system , logarithmic Sobolev inequality
2020 MSC:
35J91 , 35J20 , 35B38 , 35Qxx
1 Introduction and main results
The study of ground state problems for the Laplacian operator was mainly motivated by the search of stationary states for some nonlinear evolution equations (e.g. of the Klein-Gordon or Schrödinger type), and some other applications in Physics and Biology (e.g. in statistical mechanics [4], cosmology [11], or population dynamics [5, 12]).
Given a potential $G:\mathbb{R}^{m}\to\mathbb{R}$ having a critical point at 0, and satisfying some growth conditions near 0 as well as asymptotically, we consider the semilinear system
$-\Delta u(x)=\nabla G(u(x)),\ u:\mathbb{R}^{d}\to\mathbb{R}^{m},\,d\geq 2,\,m\geq 1,$ (1)
whose associate energy functional is $S(u)=\int_{\mathbb{R}^{d}}\big[\frac{1}{2}|\nabla u|^{2}-G(u)\big]$. The solutions of interest are those converging asymptotically to 0, and minimizing $S$ among all nonzero solutions in an appropriate class. Such a solution is called a ground state. Among the main contributions to solve this problem, we have to mention the seminal paper of Berestycki and Lions [7], which was completed by the work of Berestycki, Gallouët, and Kavian [6], in the two dimensional limiting case, and by the work of Brezis and Lieb [8] in the vector case. The method developed in all these papers is based on a constrained minimization. For more recent developments on these results we refer for instance to the papers [3, 10, 14, 15]. It is also worth noting that in view of the Sobolev embeddings, the case of dimensions $d\geq 3$ differs from the two dimensional limiting case.
Due to the increasing interest for higher order models and their applications (e.g. in nonlinear elasticity [1] and mechanics [17]), it is also relevant to investigate ground state problems for the bilaplacian operator $\Delta^{2}$. This was done by Mederski and Siemienowski [16] for dimensions $d\geq 5$, and as an application they conceived a new biharmonic logarithmic Sobolev inequality (cf. [16, Theorem 1.4]).
The scope of the present paper is to study the limiting case of space dimension $4$. Our approach is based on the method introduced in [8], as well as on some tools borrowed from [16] (e.g. Pohozaev identity in [16, Theorem 1.2]). Following [16], we also extend the biharmonic logarithmic Sobolev inequality in [16, Theorem 1.4] to dimension $4$.
In this context, we consider the system
$\Delta^{2}u(x)=g(u(x)),\ u:\mathbb{R}^{4}\to\mathbb{R}^{m},\,m\geq 1,$ (2)
where the map $g:\mathbb{R}^{m}\to\mathbb{R}^{m}$, $g\in L^{\infty}_{loc}(\mathbb{R}^{m})$ [^1] To simplify the notation, we write $g\in L^{\infty}_{loc}(\mathbb{R}^{m})$ instead of $g\in L_{loc}^{\infty}(\mathbb{R}^{m};\mathbb{R}^{m})$, and when considering other vector-valued maps belonging to some $L^{p}$ space, we shall in general omit the range., is the gradient (except at $u=0$) of a function $G:\mathbb{R}^{m}\to\mathbb{R}$ satisfying the following assumptions:
$G(0)=0,\ G:\mathbb{R}^{m}\to\mathbb{R}\text{ is continuous, and }G\in C^{1}(\mathbb{R}^{4}\setminus\{0\}),$ (3)
$g(0)=0,\text{ and }g(u)=\nabla G(u),\,\forall u\neq 0,$ (4)
$G(u)<0\ \text{ for }0<|u|\leq\epsilon,\text{ and for some }\epsilon>0,$ (5)
$G(u_{0})>0\ \text{for some}\ u_{0}.$ (6)
That is, the function $G$ has at $u=0$ a local (but not global) maximum equal to 0. We also point out that $G$ need not be differentiable at 0. For instance, we may have G(u)=|u|G(u)=-|u| near 0. The energy functional associated with system (2) is
$S(u)=K(u)-V(u),$ (7)
$\text{where }K(u)=\dfrac{1}{2}\int_{\mathbb{R}^{4}}\left|D^{2}u(x)\right|^{2}dx=\dfrac{1}{2}\sum_{i,j=1}^{4}\int_{\mathbb{R}^{4}}\Big|\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\Big|^{2}dx,$ (8)
$\text{ and }V(u)=\int_{\mathbb{R}^{4}}G(u(x))dx.$ (9)
Our goal is to show that $S(u)$ admits a minimum in the set of nontrivial solutions $u\neq 0$ to system (2), converging to 0 as $|x|\to\infty$, in some weak sense. Such a least energy solution is called a ground state. Following [8], we shall look for this solution in the class:
$\mathcal{C}=\left\{u\in L^{1}_{loc}(\mathbb{R}^{4})\ \middle|\begin{aligned} &\ \nabla u\in D^{1,2}(\mathbb{R}^{4}),\ G(u)\in L^{1}(\mathbb{R}^{4})\\ &\ \mu(\{|u|>\eta\})<\infty,\ \forall\eta>0\end{aligned}\right\},$
where $\mu$ denotes the Lebesgue measure, and by definition of the space $D^{1,2}$ (cf. [20]), we have
$\nabla u\in D^{1,2}(\mathbb{R}^{4})\Leftrightarrow\nabla u\in L^{4}(\mathbb{R}^{4}),\ D^{2}u\in L^{2}(\mathbb{R}^{4}).$
It is worth noting that $K(u)=\frac{1}{2}\int_{\mathbb{R}^{4}}|\Delta u|^{2}$, provided that $\nabla u\in D^{1,2}(\mathbb{R}^{4})$ (cf. Lemma 11). Thus, $D^{2}u$ is interchangeable with $\Delta u$ in the expression of functional $K$. We shall also see in Lemma 1 below, that the class $\mathcal{C}$ is contained in $L^{\infty}(\mathbb{R}^{4})$. Due to this property, and in contrast with the case of the second order system (1), none asymptotic condition on $G$ is required to prove the existence of a ground state solution to (2).
Theorem 1.1.
If assumptions (3)-(6) hold, then
$T=\inf\left\{\int_{\mathbb{R}^{4}}\dfrac{1}{2}|D^{2}u|^{2}:\ u\in\mathcal{C}\ ,\ \int_{\mathbb{R}^{4}}G(u)\geq 0\ ,\ u\neq 0\right\}$ (10)
is achieved by some $\bar{u}\in\mathcal{C}$, with $\bar{u}\neq 0$ and $\int_{\mathbb{R}^{4}}G(\bar{u})=0$, solving (2) in $\mathcal{D}^{\prime}$. Moreover, we have
  • 1.
    $\bar{u}\in C^{3,\alpha}_{loc}\cap W^{4,q}_{loc}$ for any $0<\alpha<1$ and $1\leq q<\infty$,
  • 2.
    $\lim_{|x|\to\infty}\bar{u}(x)=0$,
  • 3.
    and the solution $\bar{u}$ is a ground state, since there holds $0<S(\bar{u})\leq S(v)$ for all $v\in\mathcal{C}$ with $v\neq 0$ which are solutions of (2) in $\mathcal{D}^{\prime}$.
Remark 1.
If $g$ is locally Hölder continuous, then Theorem 1.1 provides a classical solution $\bar{u}\in C^{4}(\mathbb{R}^{4})$.
As an application of Theorem 1.1, we choose the potential
$G(u)=\begin{cases}|u|^{2}\ln|u|&\text{ for }u\neq 0\\ 0&\text{ for }u=0,\end{cases}$
which clearly satisfies assumptions (3)-(6). Moreover, for this choice of $G$, one can see that $\mathcal{C}\subset L^{2}(\mathbb{R}^{4})$. Following [16], we derive in (12) below the biharmonic logarithmic Sobolev inequality in dimension $4$ (cf. [16, Theorem 1.4] for dimensions $d\geq 5$). It is a generalization for the bilaplacian operator of the classical logarithmic Sobolev inequality
$\ln\left(\dfrac{1}{2\pi e}\int_{\mathbb{R}^{4}}|\nabla u|^{2}\right)\geq\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|,\text{ for }u\in H^{1}(\mathbb{R}^{4}),\int_{\mathbb{R}^{4}}|u|^{2}=1,$ (11)
given in [19], which is optimal in the scalar case (cf. [9]).
Theorem 1.2.
Let $\mathcal{C}^{\prime}=\{u\in\mathcal{C}:\int_{\mathbb{R}^{4}}|u|^{2}=1\}$. For any $u\in\mathcal{C}^{\prime}$ there holds
$\dfrac{1}{2}\ln\left(\dfrac{1}{2T}\int_{\mathbb{R}^{4}}|\Delta u|^{2}\right)=\dfrac{1}{2}\ln\left(\dfrac{1}{2T}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\right)\geq\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|,$ (12)
with $\frac{1}{2T}<\frac{1}{(\pi e)^{2}}$, where $T>0$ is the minimum in (10). Moreover, the equality in (12) holds provided that $u=\bar{u}/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$ and $\bar{u}$ is a ground state solution to (2). Conversely, if the equality in (12) holds for $u\in\mathcal{C}^{\prime}$, then there are uniquely determined $\mu>0$ and $r>0$, such that $\bar{u}(x)=\mu u(rx)$ is a ground state solution to (2).
Remark 2.
Equality case in (12) is achieved for the normalized ground state solutions provided by Theorem 1.1. However, these solutions are not known explicitly, thus we neither know the exact value of the optimal constant in (12). By comparing it with the optimal constant (cf. [9]) of the classical logarithmic Sobolev inequality (11), we can at least obtain that $\frac{1}{2T}<\frac{1}{(\pi e)^{2}}$.
The proofs of Theorems 1.1 and 1.2 are given in Sections 2 and 3 respectively, while in the Appendix we gather some technical lemmas.
2 Proof of Theorem 1.1
Let $\{u_{j}\}\subset\mathcal{C}$ be a minimizing sequence for (10), i.e.
$\lim_{j\to\infty}\dfrac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}u_{j}|^{2}=T.$
Since in the limiting case of dimension $4$, $K(u)$ is invariant under scaling, we may assume that
$\mu(\{|u_{j}|>\varepsilon\})=1.$ (13)
It is clear that the norms $\|D^{2}u_{j}\|_{L^{2}(\mathbb{R}^{4})}$ are uniformly bounded. Therefore, due to the Sobolev imbedding of $D^{1,2}(\mathbb{R}^{4})$ in $L^{4}(\mathbb{R}^{4})$ (cf. [20, Theorem 1.8]), we also obtain that the norms $\|\nabla u_{j}\|_{L^{4}(\mathbb{R}^{4})}$ are uniformly bounded.
Next, we establish some properties of the class $\mathcal{C}$, and the sequence $\{u_{j}\}$.
Lemma 1.
Let $v\in\mathcal{C}$, then $v\in L^{\infty}(\mathbb{R}^{4})\cap C^{\alpha}(\mathbb{R}^{4})$, $\forall\alpha\in(0,1)$, and it holds $\lim_{|x|\to\infty}v(x)=0$. In addition, the sequence $\{u_{j}\}$ is equibounded as well as equicontinuous in $\mathbb{R}^{4}$, and satisfies
$\int_{\{|u_{j}|>\varepsilon\}}|u_{j}|^{q}\leq C_{q},\ \forall q\in[1,+\infty),\ \forall j\text{ and for some constant }C_{q}.$ (14)
Proof.
We have for every $q\geq 1$ the following estimates.
$\|\psi\|_{L^{q}(\mathbb{R}^{4})}\leq C^{\prime}_{q}\|\nabla\psi\|_{L^{2}(\mathbb{R}^{4})}(\mu(supp\psi))^{1/q},$ (15a)
holding for $\psi\in L^{1}_{loc}(\mathbb{R}^{4})$ such that $\nabla\psi\in L^{2}(\mathbb{R}^{4})$ and $\mu(supp\psi)<\infty$,
$\|\phi\|_{L^{q}(\mathbb{R}^{4})}\leq C^{\prime}_{q}\|\nabla\phi\|_{L^{4}(\mathbb{R}^{4})}(\mu(supp\phi))^{\frac{1}{q}+\frac{1}{4}},$ (15b)
holding for $\phi\in L^{1}_{loc}(\mathbb{R}^{4})$, such that $\nabla\phi\in L^{4}(\mathbb{R}^{4})$ and $\mu(supp\phi)<\infty$, where $C^{\prime}_{q}>0$ is a constant. Estimate (15a) can be found in [8, (3.10)], while (15b) follows from (15a) and the Cauchy-Schwarz inequality, since we have $\nabla\phi\in L^{2}(\mathbb{R}^{4})$ and $\|\nabla\phi\|_{L^{2}(\mathbb{R}^{4})}\leq\|\nabla\phi\|_{L^{4}(\mathbb{R}^{4})}(\mu(supp\phi))^{1/4}$.
To see that any $v\in\mathcal{C}$ is bounded, we use the fact that $\nabla v\in L^{4}(\mathbb{R}^{4})$, which implies that $\mu(\{|\nabla v|>1\})<\infty$. Then, applying (15a) to the function $\psi=(|\nabla v|-1)^{+}$, we obtain that
$\gamma_{q}=\int_{|\nabla v|>1}|\nabla v|^{q}$ $\leq\int_{|\nabla v|>1}((|\nabla v|-1)^{+}+1)^{q}$
$\leq(\|\psi\|_{L^{q}(\mathbb{R}^{4})}+\mu(\{|\nabla v|>1\}))^{q}<\infty,$
for any $q\in[1,\infty)$. Consequently, it follows that the integrals
$\int_{B(x,1)}|\nabla v|^{q}$ $=\int_{\{|\nabla v|>1\}\cap B(x,1)}|\nabla v|^{q}+\int_{\{|\nabla v|\leq 1\}\cap B(x,1)}|\nabla v|^{q}$
$\leq\gamma_{q}+\mu(B(0,1)),$
are uniformly bounded on the unit balls of $\mathbb{R}^{4}$. A similar application of (15b) to $\phi=(|v|-1)^{+}$ shows that the integrals $\int_{B(x,1)}|v|^{q}$ are also uniformly bounded. Therefore the $W^{1,q}$ norm of $v$ is uniformly bounded on the unit balls, for any $q\in[1,\infty)$. This implies that $v\in L^{\infty}(\mathbb{R}^{4})\cap C^{\alpha}(\mathbb{R}^{4})$, for any $\alpha\in(0,1)$.
A consequence of the uniform continuity of $v$ is that the set $\{|v|>\epsilon\}$ is bounded. Indeed, if $|v(x_{n})|>\epsilon$ holds for a sequence $\{x_{n}\}\subset\mathbb{R}^{4}$ such that $\lim_{n\to\infty}|x_{n}|=\infty$, then there exists $\delta>0$ such that $|v|>\epsilon/2$ holds on the balls $B(x_{n},\delta)$, which by passing to a subsequence can be assumed to be disjoint. Consequently, we obtain that $\mu(\{|v|>\epsilon/2\})=\infty$, a contradiction.
Now, let us assume the existence of a sequence $\{x_{n}\}\subset\mathbb{R}^{4}$ satisfying $\lim_{n\to\infty}|x_{n}|=\infty$, and $|v(x_{n})|>\eta$, for some $\eta\in(0,\epsilon)$. According to what precedes, we have $|v(x_{n})|\in[\eta,\epsilon]$ (for $n$ large enough). Then, in view of the uniform continuity of the function $\mathbb{R}^{4}\ni x\mapsto G(v(x))$ and (5), there exists $\delta>0$, such that $|G(v)|>\beta>0$ holds on the balls $B(x_{n},\delta)$, for some $\beta>0$. As previously, the balls $B(x_{n},\delta)$ can be assumed to be disjoint, thus we get $G(v)\notin L^{1}(\mathbb{R}^{4})$, a contradiction. Therefore, we established that $\lim_{|x|\to\infty}v(x)=0$.
Finally, since the measures $\mu(\{|\nabla u_{j}|>1\})$ and $\mu(\{|u_{j}|>\varepsilon\})$ are uniformly bounded, the same arguments applied to the sequences $\psi_{j}=(|\nabla u_{j}|-1)^{+}$ and $\phi_{j}=(|u_{j}|-\epsilon)^{+}$ imply (14), and the fact that the $W^{1,q}$ norms of $\{u_{j}\}$ are uniformly bounded on the unit balls, for any $q\in[1,\infty)$. Thus, the sequence $\{u_{j}\}$ is equibounded and equicontinuous in $\mathbb{R}^{4}$. ∎
At this stage, we recall the following compactness result [13, Lemma 6].
Lemma 2.
Let $v\in L^{4}_{loc}(\mathbb{R}^{4})$, $\nabla v\in L^{4}(\mathbb{R}^{4})$, $\|\nabla v\|_{L^{4}}\leq C$ and $\mu(\{|v|>\varepsilon\})\geq\delta>0$. Then, there exists a shift $T_{y}v(x)=v(x+y)$ such that, for some constant $\beta=\beta(C,\delta,\varepsilon)>0$, $\mu(B\cap\{|T_{y}v|>\varepsilon/2\})>\beta$, where $B=\{x\in\mathbb{R}^{4}:|x|\leq 1\}$.
In view of Lemma 2, we can shift each $u_{j}$ in such a way that $\mu(B\cap\{|T_{x_{j}}u_{j}|>\epsilon/2\})\geq\beta$. Thus, we may assume that $\mu(B\cap\{|u_{j}|>\epsilon/2\})\geq\beta$. On the other hand, the previous estimates imply that we can extract a subsequence still called $\{u_{j}\}$ such that
  1. 1.
    $u_{j}\to u$ uniformly on compact subsets of $\mathbb{R}^{4}$,
  2. 2.
    $\nabla u_{j}\rightharpoonup\nabla u$ in $L^{4}(\mathbb{R}^{4})$,
  3. 3.
    $D^{2}u_{j}\rightharpoonup D^{2}u$ in $L^{2}(\mathbb{R}^{4})$,
  4. 4.
    $\mu(B\cap\{|u|\geq\varepsilon/2\})\geq\beta>0\ \Rightarrow u\not\equiv 0$,
  5. 5.
    $\mu(\{|u|>\varepsilon\})\leq 1$,
  6. 6.
    $u\in\mathcal{C}$.
Properties 4. and 5. of the limit $u$ follow from the convergence in 1., which is established via the theorem of Ascoli and a diagonal argument. To check that $G(u)\in L^{1}$, we use assumption (5) and the inequality $G^{+}(v)\leq k|v|$, holding for $|v|\leq\sup_{j}\|u_{j}\|_{L^{\infty}(\mathbb{R}^{4})}$, and for a constant $k>0$. Then, (14) implies that the integrals $\int_{\mathbb{R}^{4}}G^{+}(u_{j})$ are uniformly bounded, and we deduce from Fatou’s lemma that $\int_{\mathbb{R}^{4}}G^{+}(u)<\infty$. Similarly, since we have $\int_{\mathbb{R}^{4}}G^{-}(u_{j})\leq\int_{\mathbb{R}^{4}}G^{+}(u_{j})$, the integrals $\int_{\mathbb{R}^{4}}G^{-}(u_{j})$ are uniformly bounded, and we obtain as previously that $\int_{\mathbb{R}^{4}}G^{-}(u)<\infty$. Therefore, $G(u)\in L^{1}$. To prove that $u\in\mathcal{C}$, it remains to show that $\mu(\{|u|>\eta\})<\infty$, $\forall\eta>0$. Indeed, if $\eta\geq\epsilon$, this is clear in view of 5. above. Otherwise, if $0<\eta<\epsilon$, we have
$\mu(\{\eta<|u|\leq\epsilon\})\min\{|G(v)|:\eta\leq|v|\leq\epsilon\}\leq\int_{\eta\leq|u|\leq\epsilon}|G(u)|<\infty,$
thus $\mu(\{\eta<|u|\leq\epsilon\})<\infty\Rightarrow\mu(\{|u|>\eta\})<\infty$.
Following [8], we also define the class
$\mathcal{K}=\left\{\phi:\mathbb{R}^{4}\to\mathbb{R}^{m}|\ \phi\in C_{0}(\mathbb{R}^{4}),\ \nabla\phi\in L^{4},\ D^{2}\phi\in L^{2}\right\},$ (16)
where $C_{0}$ is the space of continuous functions with compact support.
In the sequel, we proceed in several steps as in [8], to construct the solution of Theorem 1.1.
Lemma 3.
Let $\phi\in\mathcal{K}$ be such that $\int_{\mathbb{R}^{4}}(G(u+\phi)-G(u))>0$. Then,
$\int_{\mathbb{R}^{4}}D^{2}u\cdot D^{2}\phi+\dfrac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}\phi|^{2}\geq 0,$ (17)
where $D^{2}u\cdot D^{2}\phi=\sum_{i,j=1}^{4}\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\cdot\frac{\partial^{2}\phi}{\partial x_{i}\partial x_{j}}$.
Proof.
Let $\phi\in\mathcal{K}$ be such that $\int_{\mathbb{R}^{4}}(G(u+\phi)-G(u))>0$. Due to the uniform convergence $G(u_{j}+\phi)-G(u_{j})\to G(u+\phi)-G(u)$ on $supp\phi$, we have that
$\int_{\mathbb{R}^{4}}G(u_{j}+\phi)$ $=\int_{supp\phi}G(u_{j}+\phi)+\int_{\mathbb{R}^{4}\setminus supp\phi}G(u_{j})$
$\geq\int_{supp\phi}(G(u_{j}+\phi)-G(u_{j}))>0,$
for $j$ large enough. Thus,
$\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}(u_{j}+\phi)|^{2}=\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}(u_{j})|^{2}+\int_{\mathbb{R}^{4}}D^{2}u_{j}\cdot D^{2}\phi+\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}\phi|^{2}\geq T.$ (18)
Since, $\lim_{j\to\infty}\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}(u_{j})|^{2}=T$, by letting $j\to\infty$ in (18), we get (17). ∎
Lemma 4.
There exists a constant $C_{G}$ (depending only on $G$) such that if $\phi\in\mathcal{K}$, then
$\int_{\mathbb{R}^{4}}g(u)\cdot\phi-C_{G}\int_{u=0}|\phi|>0\ \Rightarrow\ \int_{\mathbb{R}^{4}}D^{2}u\cdot D^{2}\phi\geq 0.$ (19)
Proof.
Let $\phi\in\mathcal{K}$. By dominated convergence, one can see that as $t\to 0^{+}$, we have
$\int_{u\neq 0}\big(G(u+t\phi)-G(u)\big)=t\int_{u\neq 0}g(u)\cdot\phi+o(t)=t\int_{\mathbb{R}^{4}}g(u)\cdot\phi+o(t).$ (20)
On the other hand, it is clear that
$\left|\int_{u=0}G(t\phi)\right|\leq C_{G}t\int_{u=0}|\phi|,\text{ as }t\to 0^{+},$ (21)
holds for a constant $C_{G}>0$ depending only on $G$, since we have the bound $|G(v)|\leq C_{G}|v|$, provided that $|v|\leq\epsilon$.
Now, choose $\phi\in\mathcal{K}$ such that $\int_{\mathbb{R}^{4}}g(u)\cdot\phi-C_{G}\int_{u=0}|\phi|>0$. Then, for small enough $t>0$, it follows from (20) and (21) that $\int_{\mathbb{R}^{4}}\big(G(u+t\phi)-G(u)\big)>0$. By Lemma 3, we conclude that
$\int_{\mathbb{R}^{4}}D^{2}u\cdot D^{2}\phi+\dfrac{t}{2}\int_{\mathbb{R}^{4}}|D^{2}\phi|^{2}\geq 0,\ \text{ for }0<t\ll 1\Rightarrow\int_{\mathbb{R}^{4}}D^{2}u\cdot D^{2}\phi\geq 0.$
Lemma 5.
Consider the linear functional $L(\phi)=\int_{\mathbb{R}^{4}}D^{2}u\cdot D^{2}\phi$. There is some $\phi_{0}\in\mathcal{K}$ such that $L(\phi_{0})\neq 0$ and $\phi_{0}=0$ on the set $\{u=0\}$.
Proof.
Let $\phi_{n}=u\rho_{n}\in\mathcal{K}$, where $\rho_{n}=\rho(x/n)$, and $\rho\in C^{\infty}_{0}(\mathbb{R}^{4})$ is a function such that $0\leq\rho\leq 1$ and
$\rho(x)=\begin{cases}1&\text{ for }|x|\leq 1\\ 0&\text{ for }|x|\geq 2.\end{cases}$ (22)
Let also $A_{n}=\{x\in\mathbb{R}^{4}:n\leq|x|\leq 2n\}$. We assume by contradiction that $L(\phi)=0$, for all $\phi\in\mathcal{K}$ such that $\phi=0$ on $\{u=0\}$. It is clear that $\phi_{n}=0$ on the set $\{u=0\}$. Thus, we have $L(\phi_{n})=L(u\rho_{n})=0$. That is,
$\int_{\mathbb{R}^{4}}\sum_{i,j=1}^{4}\Big(\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\cdot\Big(\frac{\partial u}{\partial x_{i}}\frac{\partial\rho_{n}}{\partial x_{j}}+\frac{\partial u}{\partial x_{j}}\frac{\partial\rho_{n}}{\partial x_{i}}\Big)+\Big(u\cdot\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\Big)\frac{\partial^{2}\rho_{n}}{\partial x_{i}\partial x_{j}}\Big)\\ +\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\rho_{n}=0.$
After direct calculations we obtain that
$\lim_{n\to\infty}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\rho_{n}=\int_{\mathbb{R}^{4}}|D^{2}u|^{2},$
$\lim_{n\to\infty}\int_{\mathbb{R}^{4}}\Big|\Big(\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\cdot\frac{\partial u}{\partial x_{i}}\Big)\frac{\partial\rho_{n}}{\partial x_{j}}\Big|\leq k_{\rho}\|D^{2}u\|_{L^{2}(A_{n})}\|\nabla u\|_{L^{4}(A_{n})}\to 0\text{, as }n\to\infty,$
$\lim_{n\to\infty}\int_{\mathbb{R}^{4}}\Big|\Big(u\cdot\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\Big)\frac{\partial^{2}\rho_{n}}{\partial x_{i}\partial x_{j}}\Big|\leq k_{\rho}\|D^{2}u\|_{L^{2}(A_{n})}\|u\|_{L^{\infty}(\mathbb{R}^{4})}\to 0\text{, as }n\to\infty,$
for a constant $k_{\rho}>0$ depending on $\rho$. This implies that $D^{2}u\equiv 0$, and since $u$ is bounded, we get $u\equiv 0$, which is a contradiction. ∎
Lemmas 3.5 and 3.6 from [8] are applicable to our case without modifications. For convenience of the reader we reproduce in Lemmas 6 and 7 below their proofs.
Lemma 6.
There is a constant $\lambda\geq 0$ such that
$\Big|\int_{\mathbb{R}^{4}}g(u)\cdot\phi-\lambda L(\phi)\Big|\leq C_{G}\int_{u=0}|\phi|,\ \forall\phi\in\mathcal{K}.$ (23)
Proof.
In view of Lemma 5, we can choose $\phi_{0}\in\mathcal{K}$ such that $\phi_{0}=0$ on $\{u=0\}$, and $L(\phi_{0})=-1$. Next, given $\phi\in\mathcal{K}$, we consider the function $\psi=\phi+L(\phi)\phi_{0}+\alpha\phi_{0}$, with $\alpha>0$. It is clear that $L(\psi)=-\alpha<0$, thus by Lemma 4, we have that $\int_{\mathbb{R}^{4}}g(u)\cdot(\phi+L(\phi)\phi_{0}+\alpha\phi_{0})-C_{G}\int_{u=0}|\phi|\leq 0.$ Setting $\lambda=-\int_{\mathbb{R}^{4}}g(u)\cdot\phi_{0}\geq 0$ (by (19)), and letting $\alpha\to 0$, we deduce that $\int_{\mathbb{R}^{4}}g(u)\cdot\phi-\lambda L(\phi)-C_{G}\int_{u=0}|\phi|\leq 0$, $\forall\phi\in\mathcal{K}$, from which (23) follows. ∎
Lemma 7.
We have that $\frac{G(u)}{\partial x_{i}}=g(u)\cdot\frac{\partial u}{\partial x_{i}}$ in $\mathcal{D}^{\prime}$.
Proof.
Let $M=\|u\|_{L^{\infty}(\mathbb{R}^{4})}$, and let $G_{k}$ be a smoothing sequence for $G$ so that $G_{k}\to G$ uniformly on $B_{M}=\{v\in\mathbb{R}^{m}:|v|\leq M\}$, and $g_{k}=\nabla G_{k}\to g$ pointwise on $B_{M}\setminus\{0\}$. Moreover, we can ensure that the maps $g_{k}$ are uniformly bounded on $B_{M}$. It is clear that $\frac{\partial G_{k}(u)}{\partial x_{i}}=g_{k}(u)\cdot\frac{\partial u}{\partial x_{i}}$. Thus, given $\phi\in C^{\infty}_{0}$ we have $\int_{\mathbb{R}^{4}}\frac{\partial\phi}{\partial x_{i}}G_{k}(u)=-\int_{\mathbb{R}^{4}}\phi\big(g_{k}(u)\cdot\frac{\partial u}{\partial x_{i}}\big)$, and passing to the limit we get $\int_{\mathbb{R}^{4}}\frac{\partial\phi}{\partial x_{i}}G(u)=-\int_{\mathbb{R}^{4}}\phi\big(g(u)\cdot\frac{\partial u}{\partial x_{i}}\big)$, since $\frac{\partial u}{\partial x_{i}}=0$ on the set $\{u=0\}$. ∎
Lemma 8.
Let $\lambda$ be as in Lemma 6. Then, $u$ is a solution of the equation $\Delta^{2}u=\frac{1}{\lambda}g(u)$ in $W^{4,q}_{loc}$, $\forall q\in[1,\infty)$, satisfying $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}=T$ and $\int_{\mathbb{R}^{4}}G(u)=0$.
Proof.
The linear functional $M(\phi)=\int_{\mathbb{R}^{4}}g(u)\cdot\phi-\lambda L(\phi)$ restricted to $C^{\infty}_{0}$ satisfies $|M(\phi)|\leq C_{G}\|\phi\|_{L^{1}}$ (cf. (23)), thus by the Riesz Representation theorem, there exists $h\in L^{\infty}(\mathbb{R}^{4})$, $h:\mathbb{R}^{4}\to\mathbb{R}^{m}$, such that $M(\phi)=\int_{\mathbb{R}^{4}}h\cdot\phi$, $\forall\phi\in C^{\infty}_{0}$. Actually, (23) implies that $\int_{\mathbb{R}^{4}}h\cdot\phi=0$, for all $\phi\in C^{\infty}_{0}$ with support in the open set $\{u\neq 0\}$. Consequently, $h$ vanishes on $\{u\neq 0\}$, and denoting by $\chi$ the characteristic function, we have $\lambda\Delta^{2}u+g(u)=h\chi_{u=0}$ in $\mathcal{D}^{\prime}$. Our next claim is that $\lambda>0$. Otherwise, if $\lambda=0$, then $g(u)\equiv 0$. Moreover, in view of Lemma 7, the continuous function $x\mapsto G(u(x))$ is constant. Since $\lim_{|x|\to\infty}G(u(x))=0$, it follows that $G(u)\equiv 0$ and $u\equiv 0$, which is a contradiction. Hence, it holds $\Delta^{2}u=\frac{1}{\lambda}(g(u)-h\chi_{u=0})$ in $\mathcal{D}^{\prime}$, with $\lambda>0$. Finally, since $u\in L^{\infty}$, while $g(u)-h\chi_{u=0}\in L^{\infty}$, we know by Lemma 9 that $u\in W^{4,q}_{loc}$, $\forall q\geq 1$. Therefore, $\Delta^{2}u=0$ a.e. on $\{u=0\}$, and $h=0$ a.e. on $\{u=0\}$. That is, we have $\Delta^{2}u=\frac{1}{\lambda}g(u)$ in $\mathcal{D}^{\prime}$, with $\lambda>0$. In addition, Pohozaev identity (cf. Lemma 10) implies that $\int_{\mathbb{R}^{4}}G(u)=0$, while in view of the weak convergence $D^{2}u_{j}\rightharpoonup D^{2}u$ in $L^{2}(\mathbb{R}^{4})$, we have $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\leq T$. Thus, $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}=T$, and $u$ solves the minimization problem (10). ∎
Setting $\bar{u}(x)=u(\lambda^{-\frac{1}{4}}x)$, we obtain a solution of (2), such that $\int_{\mathbb{R}^{4}}G(\bar{u})=0$, and $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}\bar{u}|^{2}=T$ (i.e. $\bar{u}$ still solves the minimization problem (10)). Next, let $v\in\mathcal{C}$ be a solution of (2) in $\mathcal{D}^{\prime}$ such that $v\neq 0$. Then, by Pohozaev identity it holds $\int_{\mathbb{R}^{4}}G(v)=0$, and we have $S(v)=K(v)\geq T=K(\bar{u})=S(\bar{u})$. That is, $\bar{u}$ has all the properties stated in Theorem 1.1. We also point out that if $g$ is locally Hölder continuous, then by elliptic regularity we obtain that $\bar{u}\in C^{4}(\mathbb{R}^{4})$, since the function $x\mapsto g(\bar{u}(x))$ is locally Hölder continuous.
Remark 3.
Any solution $u$ of the minimization problem (10) solves the equation $\Delta^{2}u=\frac{1}{\lambda}g(u)$ in $\mathcal{D}^{\prime}$, where $\lambda>0$ is the Lagrange multiplier provided by Lemma 6. In addition, there holds $\int_{\mathbb{R}^{4}}G(u)=0$. Indeed, by considering the minimizing sequence $u_{j}=u$, $\forall j$, the arguments of Lemmas 3, 4, 5, 6, 7 and 8 can be reproduced and apply to the minimizer $u$. Thus, the map $\bar{u}(x)=u(rx)$, with $r=\lambda^{-\frac{1}{4}}$, is a ground state solution to (2).
3 Proof of Theorem 1.2
As we mentioned in the Introduction, given the potential $G(u)=|u|^{2}\ln|u|$ (which clearly satisfies assumptions (3)-(6)), one can see that $\mathcal{C}\subset L^{2}(\mathbb{R}^{4})$. Let $u\in\mathcal{C}^{\prime}$, and set $\mu=e^{-V(u)}$, with $V(u)=\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|$. We compute $V(\mu u)=\mu^{2}(\ln(\mu)\|u\|^{2}_{L^{2}(\mathbb{R}^{4})}+V(u))=0$, since $\|u\|^{2}_{L^{2}(\mathbb{R}^{4})}=1$. Consequently, we have the inequality $\frac{\mu^{2}}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\geq T$, which is equivalent to (12). Next, assume that $u=\bar{u}/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$, where $\bar{u}$ is a ground state solution to (2). Then, setting $\mu=1/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$ one can see that the right hand side of (12) is equal to $\frac{1}{2}\ln(\mu^{2})$, while the left hand side is $V(\mu\bar{u})=\mu^{2}(\ln(\mu)\|\bar{u}\|^{2}_{L^{2}(\mathbb{R}^{4})}+V(\bar{u}))=\ln(\mu)$, since $V(\bar{u})=0$. This is an equality case in (12). Conversely, if the equality in (12) holds for $u\in\mathcal{C}^{\prime}$, i.e. $K(u)=Te^{2V(u)}$ (where $K$ is defined in (7)), then we have $K(\mu u)=T\Leftrightarrow\mu=e^{-V(u)}$. Moreover, for this value of $\mu$ there holds $V(\mu u)=\mu^{2}(\ln(\mu)\|u\|^{2}_{L^{2}(\mathbb{R}^{4})}+V(u))=0$, that is, $\mu u$ solves the minimization problem (10). Therefore, in view of Remark 3, there exists a unique $r>0$ such that $\bar{u}(x)=\mu u(rx)$ is a ground state solution to (2).
To complete the proof of the theorem, it remains to show that $\frac{1}{2T}<\frac{1}{(\pi e)^{2}}$. To this end, we use (11) (which also holds in the vector case since we have $|\nabla|u||\leq|\nabla u|$, $\forall u\in H^{1}(\mathbb{R}^{4};\mathbb{R}^{m})$), combined with the inequality (cf. [16, Lemma 6.1]):
$\int_{\mathbb{R}^{4}}|\nabla u|^{2}<\Big[\int_{\mathbb{R}^{4}}|\Delta u|^{2}\Big]^{\frac{1}{2}}\Big[\int_{\mathbb{R}^{4}}|u|^{2}\Big]^{\frac{1}{2}},\text{ for }u\in H^{2}(\mathbb{R}^{4}),u\neq 0.$ (24)
In view of Lemma 11, we have $\mathcal{C}^{\prime}\subset H^{2}(\mathbb{R}^{4})$, as well as $\|D^{2}u\|_{L^{2}(\mathbb{R}^{4})}=\|\Delta u\|_{L^{2}(\mathbb{R}^{4})}$, thus we obtain
$\dfrac{1}{2}\ln\left(\dfrac{1}{(2\pi e)^{2}}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\right)>\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|,\text{ for }u\in\mathcal{C}^{\prime},$ (25)
and choosing $u=\bar{u}/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$ with $\bar{u}$ a ground state solution to (2), we conclude that
$\dfrac{1}{2}\ln\left(\dfrac{1}{(2\pi e)^{2}}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\right)>\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|=\dfrac{1}{2}\ln\left(\dfrac{1}{2T}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\right).$ (26)
This proves that $\frac{1}{2T}<\frac{1}{(\pi e)^{2}}$.
4 Some lemmas applicable to the solutions of (2)
Here we gather some lemmas applicable to the solutions of (2). We first recall a regularity result for the nonhomogeneous biharmonic equation that can be derived from the $L^{p}$ estimates in [18, (2.6)].
Lemma 9.
Let $1<p<\infty$, and $\Omega\subset\mathbb{R}^{n}$ an open set. If $u\in L^{p}_{loc}(\Omega)$ and $\Delta^{2}u\in L^{p}_{loc}(\Omega)$, then $u\in W^{4,p}_{loc}(\Omega)$.
Next, we state Pohozaev identity for solutions to (2) belonging to $\mathcal{C}$. This identity was established in [16, Theorem 1.2] for dimensions $d\geq 5$. Its proof is identical in dimension $4$.
Lemma 10.
Let $v\in\mathcal{C}$ be a solution to (2) in $\mathcal{D}^{\prime}$. Then, $\int_{\mathbb{R}^{4}}G(v)=0$.
Proof.
We notice that since $v$ as well as $g(v)$ are bounded in $\mathbb{R}^{4}$, it follows from Lemma 9 that $v\in W^{4,q}_{loc}(\mathbb{R}^{4})$, $\forall q\geq 1$. Then, we reproduce the arguments in the proof of [16, Theorem 1.2]. ∎
Finally, we establish that $D^{2}u$ is interchangeable with $\Delta u$ in the expression of functional $K$.
Lemma 11.
Let $u\in L^{1}_{loc}(\mathbb{R}^{4})$ be such that $\nabla u\in L^{4}(\mathbb{R}^{4})$, and $D^{2}u\in L^{2}(\mathbb{R}^{4})$. Then, we have $\int_{\mathbb{R}^{4}}|D^{2}u|^{2}=\int_{\mathbb{R}^{4}}|\Delta u|^{2}$. If in addition $u\in L^{2}(\mathbb{R}^{4})$, then it holds $\nabla u\in L^{2}(\mathbb{R}^{4})$, and $\int_{\mathbb{R}^{4}}|\nabla u|^{2}=-\int_{\mathbb{R}^{4}}u\cdot\Delta u$.
Proof.
(i) Let $\phi\in C^{\infty}(\mathbb{R}^{4})$ be such that $\nabla\phi\in L^{4}(\mathbb{R}^{4})$, and $D^{2}\phi\in L^{2}(\mathbb{R}^{4})$. By integrating by parts we obtain
$\int_{\mathbb{R}^{4}}|D^{2}\phi|^{2}=\lim_{r\to\infty}\sum_{i,j=1}^{4}\int_{B_{r}}\partial^{2}_{ij}\phi\cdot\partial^{2}_{ij}\phi$
$=$ $\lim_{r\to\infty}\sum_{i,j=1}^{4}\Big[-\int_{B_{r}}\partial_{i}\phi\cdot\partial^{3}_{ijj}\phi+\int_{\partial B_{r}}\nu_{j}\partial_{i}\phi\cdot\partial^{2}_{ij}\phi\Big]$
$=$ $\lim_{r\to\infty}\sum_{i,j=1}^{4}\Big[\int_{B_{r}}\partial^{2}_{ii}\phi\cdot\partial^{2}_{jj}\phi+\int_{\partial B_{r}}(\nu_{j}\partial_{i}\phi\cdot\partial^{2}_{ij}\phi-\nu_{i}\partial_{i}\phi\cdot\partial^{2}_{jj}\phi)\Big]$
$=$ $\int_{\mathbb{R}^{4}}|\Delta\phi|^{2}+\lim_{r\to\infty}\sum_{i,j=1}^{4}\Big[\int_{\partial B_{r}}(\nu_{j}\partial_{i}\phi\cdot\partial^{2}_{ij}\phi-\nu_{i}\partial_{i}\phi\cdot\partial^{2}_{jj}\phi)\Big],$
where $B_{r}=\{x\in\mathbb{R}^{4}:|x|<1\}$, and $\nu=(\nu_{1},\nu_{2},\nu_{3},\nu_{4})$ is the outward unit normal vector at $\partial B_{r}$. Moreover, since $|\nabla\phi||D^{2}\phi|\in L^{\frac{4}{3}}(\mathbb{R}^{4})$, there exists a sequence $r_{k}\to\infty$, such that $\lim_{k\to\infty}r_{k}\int_{\partial B_{r_{k}}}(|\nabla\phi||D^{2}\phi|)^{\frac{4}{3}}=0$. Therefore, we conclude that
$\Big|\int_{\partial B_{r_{k}}}(\nu_{j}\partial_{i}\phi\cdot\partial^{2}_{ij}\phi-\nu_{i}\partial_{i}\phi\cdot\partial^{2}_{jj}\phi)\Big|\leq\Big[\int_{\partial B_{r_{k}}}(|\nabla\phi||D^{2}\phi|)^{\frac{4}{3}}\Big]^{\frac{3}{4}}O(r^{\frac{3}{4}})\to 0,$
as $k\to\infty$, and $\int_{\mathbb{R}^{4}}|D^{2}\phi|^{2}=\int_{\mathbb{R}^{4}}|\Delta\phi|^{2}$. Now, if $u\in L^{1}_{loc}(\mathbb{R}^{4})$ is such that $\nabla u\in L^{4}(\mathbb{R}^{4})$, and $D^{2}u\in L^{2}(\mathbb{R}^{4})$, we consider a sequence $\rho_{n}$ of mollifiers. Setting $\phi_{n}=u\star\rho_{n}\in C^{\infty}(\mathbb{R}^{4})$, we have $\partial_{i}\phi_{n}=\partial_{i}u\star\rho_{n}\to\partial_{i}u$ in $L^{4}(\mathbb{R}^{4})$, as well as $\partial^{2}_{ij}\phi_{n}=\partial^{2}_{ij}u\star\rho_{n}\to\partial^{2}_{ij}u$ in $L^{2}(\mathbb{R}^{4})$. Consequently, it holds $\int_{\mathbb{R}^{4}}|D^{2}\phi_{n}|^{2}=\int_{\mathbb{R}^{4}}|\Delta\phi_{n}|^{2}$, and as $n\to\infty$, we get $\int_{\mathbb{R}^{4}}|D^{2}u|^{2}=\int_{\mathbb{R}^{4}}|\Delta u|^{2}$.
(ii) To prove the second statement of the lemma, we consider a function $\phi\in C^{\infty}(\mathbb{R}^{4})\cap L^{2}(\mathbb{R}^{4})$ such that $\nabla\phi\in L^{4}(\mathbb{R}^{4})$, and $D^{2}\phi\in L^{2}(\mathbb{R}^{4})$. As previously, we integrate by parts and obtain
$\int_{B_{r}}|\nabla\phi|^{2}=-\int_{B_{r}}\phi\cdot\Delta\phi+\int_{\partial B_{r}}\frac{\partial\phi}{\partial\nu}\cdot\phi.$
On the one hand, it is clear that $\lim_{r\to\infty}\int_{B_{r}}\phi\cdot\Delta\phi=\int_{\mathbb{R}^{4}}\phi\cdot\Delta\phi$. On the other hand, we notice that since $|\nabla\phi||\phi|\in L^{\frac{4}{3}}(\mathbb{R}^{4})$, there exists a sequence $r_{k}\to\infty$, such that $\lim_{k\to\infty}r_{k}\int_{\partial B_{r_{k}}}\frac{\partial\phi}{\partial\nu}\cdot\phi=0$. Therefore, we have
$\Big|\int_{\partial B_{r_{k}}}\frac{\partial\phi}{\partial\nu}\cdot\phi\Big|\leq\Big[\int_{\partial B_{r_{k}}}(|\nabla\phi||\phi|)^{\frac{4}{3}}\Big]^{\frac{3}{4}}O(r^{\frac{3}{4}})\to 0,$
as $k\to\infty$, and $\lim_{k\to\infty}\int_{B_{r_{k}}}|\nabla\phi|^{2}=-\int_{\mathbb{R}^{4}}\phi\cdot\Delta\phi$. This proves that $\nabla\phi\in L^{2}(\mathbb{R}^{4})$, and $\int_{\mathbb{R}^{4}}|\nabla\phi|^{2}=-\int_{\mathbb{R}^{4}}\phi\cdot\Delta\phi$. Finally, if $u\in L^{2}(\mathbb{R}^{4})$ is such that $\nabla u\in L^{4}(\mathbb{R}^{4})$, and $D^{2}u\in L^{2}(\mathbb{R}^{4})$, we proceed as previously by mollification. Setting $\phi_{n}=u\star\rho_{n}\in C^{\infty}(\mathbb{R}^{4})$, we have that $\phi_{n}\to u$ as well as $\Delta\phi_{n}=\Delta u\star\rho_{n}\to\Delta u$ in $L^{2}(\mathbb{R}^{4})$, while $\partial_{i}\phi_{n}=\partial_{i}u\star\rho_{n}\to\partial_{i}u$ in $L^{4}(\mathbb{R}^{4})$ and a.e. in $\mathbb{R}^{4}$ (up to subsequence). Consequently, it holds $\int_{\mathbb{R}^{4}}|\nabla\phi_{n}|^{2}=-\int_{\mathbb{R}^{4}}\phi_{n}\cdot\Delta\phi_{n}$, and as $n\to\infty$, we get that $\nabla u\in L^{2}(\mathbb{R}^{4})$, and $\int_{\mathbb{R}^{4}}|\nabla u|^{2}=-\int_{\mathbb{R}^{4}}u\cdot\Delta u$.
5 Conclusion
The method of constrained minimization applies to prove the existence of a ground state solution for both systems (1) and (2), in space dimensions greater or equal than that of the limiting case (i.e. $d\geq 2$ for (1), and $d\geq 4$ for (2)). In contrast with the second order system (1), none asymptotic condition on the potential $G$ is required for the fourth order system (2), since for the latter, the relevant class $\cal C$ of functions is contained in $L^{\infty}$. We also point out that if the space dimension is smaller than that of the limiting case, then the problem of finding nontrivial solutions converging asympototically to 0 is different. In space dimension $d=1$, such solutions of (1) are the homoclinic orbits (cf. for instance [2]). On the other hand, the construction of such solutions of (2) in space dimensions $d=1,2,3$, is to a large extent an open problem.
Acknowledgements
This research project is implemented within the framework of H.F.R.I call “Basic research Financing (Horizontal support of all Sciences)” under the National Recovery and Resilience Plan “Greece 2.0” funded by the European Union - NextGenera tionEU (H.F.R.I. Project Number: 016097).